OpenStudy (anonymous):

Please explain this to me: \[\frac{1}{2}\int_{b-1/2}^{b+1/2}\frac{1}{\sqrt{|x|}}dx=\] \[\frac{1}{2}\int_{b-1/2}^{0}\frac{1}{\sqrt{|x|}}dx+\frac{1}{2}\int_{0}^{b+1/2}\frac{1}{\sqrt{|x|}}dx=\] \[\sqrt{1/2-b}+\sqrt{1/2+b}\]

OpenStudy (anonymous):

This is the fundamental theorem of calculus. Check out this video, there's an example similar to yours at 6:00 http://www.youtube.com/watch?v=PGmVvIglZx8

OpenStudy (anonymous):

oh, also: \[0\le b \le1/2\]

OpenStudy (anonymous):

Thx, I'll take a look.

OpenStudy (anonymous):

also \[\int_a^bf(x)dx=\int_a^cf(x)dx+\int_c^bf(x)dx\] is always true

OpenStudy (anonymous):

Ok, I get the first part. What happens when the integrals are calculated?

hartnn (hartnn):

\(\int_a^bf(x)dx=\int_a^cf(x)dx+\int_c^bf(x)dx\) true if \(a \le c \le b\)

OpenStudy (anonymous):

actually it is always true, so long as the integral exists